### v1

# Convolutional Kolmogorov–Arnold Networks

Abstract
In this paper, we introduce the Convolutional Kolmogorov-Arnold Networks (Convolutional KANs), an innovative alternative to the standard Convolutional Neural Networks (CNNs) that have revolutionized the field of computer vision. We integrate the non-linear activation functions presented in Kolmogorov-Arnold Networks (KANs) into convolutions to build a new layer. Throughout the paper, we empirically validate the performance of Convolutional KANs against traditional architectures across MNIST and Fashion-MNIST benchmarks, illustrating that this new approach maintains a similar level of accuracy while using half the amount of parameters. This significant reduction of parameters opens up a new approach to advance the optimization of neural network architectures.
1 Introduction
The field of deep learning is constantly changing, the fast improvement of architectures has helped the advancement of computer vision in tasks involving complex spatial data. Convolutional Neural Networks proposed by LeCun et al.[5] are widely used due to their ability to handle high-dimensional data arrays such as images. Normally, these networks rely on linear transformations followed by an optional activation function in their convolutional layers to understand spatial relationships, which significantly reduced the number of parameters to capture complex patterns in images.
In recent years, there has been an increase in the integration of advanced mathematical theories into deep learning architectures which have helped neural networks in handling complex data structures. Kolmogorov-Arnold Networks (KANs) [6] are a promising alternative to Multi-Layer Perceptrons (MLPs)[4] that use the Kolmogorov-Arnold theorem to integrate splines which is a key component of their architecture.
In light of these advancements, this paper explores the adaptation of KANs to convolutional layers, a common element in many CNN architectures used in computer vision. Traditional CNNs utilize fixed activation functions and linear transformations which, while effective, can benefit from the flexibility and reduced parametric complexity offered by KANs. By employing spline-based convolutional layers, as proposed in SplineCNN by M. Fey and J. E. Lenssen et al. [3], networks can capture non-linear relationships more effectively.
Throughout this paper, we begin with a high-level overview of the KAN architecture to set the stage for a comprehensive mathematical treatment of Convolutional KANs. We will provide a detailed examination of different Convolutional KANs architectures and benchmark their performance against traditional models, focusing on parameter efficiency within the MNIST and Fashion-MNIST datasets. Our hypothesis posits that Convolutional KANs, by leveraging spline-based layers, will require fewer parameters while achieving accuracy levels competitive with established benchmarks, potentially setting a new standard in neural network architectures for image-related tasks [2]. For further exploration and practical application, the code for this layer is available at our GitHub repository: GitHub/Convolutional-KANs.
2 Related work
Kolmogorov-Arnold theorem and neural networks
The application of the Kolmogorov-Arnold theorem in neural networks marks a significant theoretical integration that enhances the expressiveness and efficiency of neural models. The theorem, which provides a way to represent any multivariate continuous function as a composition of univariate functions and additions, has been adapted in the design of Kolmogorov-Arnold Networks (KANs). KANs differ from traditional MLPs by replacing linear weight matrices with learnable splines, thus reducing the number of parameters required and potentially improving the generalization capabilities of the network [6].
Splines in Convolutional Neural Networks
One noteworthy development in neural network architecture involves the use of splines, particularly in the context of Convolutional Neural Networks (CNNs). The SplineCNN, as proposed by M. Fey and J. E. Lenssen et al.[3], introduces spline-based convolutional layers that enhance the network’s ability to capture non-linear relationships in the data. This approach is particularly effective in geometric deep learning, where the adaptability of splines plays a crucial role in handling non-Euclidean data.
A significant aspect of the method proposed by the authors is its treatment of images like those in the MNIST dataset, where it first processes the images by interpreting them as graphs before classification. This graph-based approach allows SplineCNN to handle irregular data structures effectively. Unlike SplineCNN, our Convolutional Kolmogorov-Arnold Networks (Convolutional KANs) apply spline functions directly on structured data such as images and matrices without necessitating their conversion into graphs.
3 Kolmogorov-Arnold Networks (KANs)
In the evolving landscape of neural networks, the use of Kolmogorov-Arnold Networks (KANs) [6] presents an innovative approach to neural network design, based on the Kolmogorov-Arnold representation theorem [1]. This theorem states that any multivariate continuous function can be represented as a composition of univariate functions and addition operations. This foundational concept sets KANs apart from traditional Multi-Layer Perceptrons (MLPs) [4].
3.1 Architecture
The core of KANs resides in their unique architecture. Unlike traditional MLPs that use fixed activation functions at nodes, KANs implement learnable activation functions on the network edges. This critical shift from static to dynamic node functions involves replacing conventional linear weight matrices with adaptive spline functions, which are parametrized and optimized during training. This allows for a more flexible and responsive model architecture that can dynamically adapt to complex data patterns.
In more detail, the Kolmogorov-Arnold representation theorem posits that a multivariate function $f(x_{1},\ldots,x_{n})$ can be expressed as:
$f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$
Here, $\phi_{q,p}$ are univariate functions mapping each input variable ($x_{p}$) such $\phi_{q,p}:[0,1]\to\mathbb{R}$, and $\Phi_{q}$ : $\mathbb{R}\to\mathbb{R}$, univariate functions.
KANs structure each layer as a matrix of these learnable 1D functions:
$\Phi=\{\phi_{q,p}\},\quad p=1,2,\ldots,n_{in},\quad q=1,2,\ldots,n_{out}$
Particularly, each function $\phi_{q,p}$ can be defined as a B-spline, a type of spline function defined by a linear combination of basis splines, enhancing the network’s ability to learn complex data representations. Here, $n_{in}$ represents the number of input features to a particular layer, while $n_{out}$ denotes the number of output features produced by that layer, reflecting the dimensionality transformations across the network layers. The activation functions $\phi_{l,j,i}$ in this matrix are such learnable spline functions, expressed as:
$\text{spline}(x)=\sum_{i}c_{i}B_{i}(x),\quad c_{i}\text{ are trainable coefficients}$
This formulation allows each $\phi_{l,j,i}$ to adapt its shape based on the data, offering unprecedented flexibility in how the network models interactions between inputs.
The overall structure of a KAN is analogous to stacking layers in MLPs, but with the enhancement of utilizing complex functional mappings instead of simple linear transformations and nonlinear activations:
$\text{KAN}(x)=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{0})(x)$
Each layer’s transformation, $\Phi_{l}$, acts on the input $x_{l}$ to produce the next layer’s input $x_{l+1}$, described as:
$x_{l+1}=\Phi_{l}(x_{l})=\left(\begin{array}[]{ccc}\phi_{l,1,1}(\cdot)&\cdots&\phi_{l,1,n_{l}}(\cdot)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\cdots&\phi_{l,n_{l+1},n_{l}}(\cdot)\end{array}\right)x_{l}$
where each activation function $\phi_{l,j,i}$ is a spline, providing a rich, adaptable response surface to model inputs.
3.2 Motivation for Using KANs
The motivation for using the architecture of KANs, with learnable activation functions on edges, enhances their expressive power and efficiency. By replacing linear weight matrices with spline functions, KANs reduce the number of parameters needed to achieve high accuracy, leading to faster convergence and better generalization.
4 Convolutional Kolmogorov-Arnold Networks
Convolutional Kolmogorov-Arnold Networks are similar to CNNs. The difference is that the Convolutional Layers are replaced by KAN Convolutional Layers and after flattening, one can either have a KAN or a MLP. The main strength of the Convolutional KANs is its requirement for significantly fewer parameters compared to other architectures. This is given by the construction of this networks, because B-Splines are able to smoothly represent aribtrary activation functions that will not be found using a ReLU in between convolutions.
4.1 KAN Convolutions
In Computer Vision, convolutions are normally used interchangeably with a mathematical operation used in Convolutional Neural-Networks. The operation consists in passing a kernel, or filter, across the input and calculating the dot product at each position. In KAN Convolutions the main idea is to propose an alternative implementation of this mathematical operation utilizing the approach of Kolmogorov-Arnold Networks. The main difference between KAN Convolutions and the convolutions used in CNNs lies in the kernel. In CNNs it is made of weights whereas in Convolutional KANs, each element of the kernel, $\phi$, is a learnable non-linear functions that utilizes B-Splines. Formally, each element is defined as:
$\phi=w_{1}\cdot spline(x)+w_{2}\cdot silu(x)\quad$ (1)
In a KAN Convolution, the kernel slides over the image and applies the corresponding activation function, $\phi_{ij}$ to the corresponding pixel, $a_{kl}$ and calculates the output pixel as the sum of $\phi_{ij}(a_{kl})$. Let K be a KAN kernel $\in\mathbb{R}^{N\times M}$, and an image as a matrix as:
$\text{Image}=\begin{bmatrix}a_{11}&a_{12}&\cdots&a_{1p}\\ a_{21}&a_{22}&\cdots&a_{2p}\\ \vdots&\vdots&\ddots&\vdots\\ a_{m1}&a_{m2}&\cdots&a_{mp}\end{bmatrix}$ (2)
Then a KAN Convolutions is defined as follows in the Equation 3
$(\text{Image}\ast K)_{i,j}=\sum_{k=1}^{N}\sum_{l=1}^{M}\phi_{kl}(a_{i+k,j+l})$ (3)
Lets see an example in matrix form:
KAN $\text{KAN Kernel}=\begin{bmatrix}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{22}\end{bmatrix}$
ImageKAN $\text{Image}\ast\text{KAN Kernel}=\begin{bmatrix}\phi_{11}(a_{11})+\phi_{12}(a_{12})+\phi_{21}(a_{21})+\phi_{22}(a_{22})&\cdots&r_{1(p-1)}\\ \phi_{11}(a_{21})+\phi_{12}(a_{22})+\phi_{21}(a_{31})+\phi_{22}(a_{32})&\cdots&r_{2(p-1)}\\ \vdots&\ddots&\vdots\\ \phi_{11}(a_{m1})+\phi_{12}(a_{m2})+\phi_{21}(a_{(m+1)1})+\phi_{22}(a_{(m+1)2})&\cdots&r_{m(p-1)}\end{bmatrix}$ (4)
4.2 Grid extension and update
The grid refers to the set of points which we want to discretize. It is initialized with a previously defined interval and control points, but during training, some of the input variables to each $\phi$ might get out of range from the grid limits. To tackle this we extended the grid to be able to capture the variables which escape the original limits. The method to extend the grid described in the original KAN paper and consists in the following optimization problem.
$\{c_{j}^{\prime}\}=\operatorname*{arg\,min}_{c^{\prime}_{j}}\mathbb{E}_{x\sim p(x)}[\sum_{j=0}^{G_{2}+k-1}c_{j}^{\prime}B(x^{\prime})_{j}^{\prime}-\sum_{j=0}^{G_{1}+k-1}c_{j}B(x)_{j}]$ (5)
Where $G_{1}$ is the previous grid size, $G_{2}$ is the new grid size and k is the B Spline degree.
While testing the models we effectively verified that the output variables of a KAN Convolutional Layer weren’t bounded to the default grid range of $[-1,1]$. This is a problem, especially when using multiple convolutional layers since the input of a convolutional layer should be in the range that the B-Spline operates in so that the "learning" is done by the splines and not the weight that modifies the SiLu. To solve this issue, during training each time an input falls outside the grid range, the grid is updated. This consists of maintaining spline shape between the original grid size and maintaining the same amount of control points, and extending the spline to a range that contains the input. Another solution for this issue is the implementation of batch normalization, after each convolutional layer a batch normalization layer is applied. This approach adds very few learneable $\mu$ and $\sigma$ parameters that Standardize the inputs to the layer to $\mu=0$ and $\sigma=1$. This ensures that most, but not all, of the outputs are between in range. As is seen in Figure 1, when the input is out of the Spline range, the layer will act as as SiLu activation as is expected, so if the range is not updated and most of inputs land out of range, a KAN will not differ from an MLP that uses SiLu activations.
a Refer to caption
Figure 1: Splines learned by the first convolution at the first position for different ranges. The left plot shows the spline learned within the range $[-1,1]$, while the right plot shows the spline learned within the range $[-10,10]$. The SILU (Sigmoid Linear Unit) function is added to the spline across the entire range, but the spline is only defined within $[-1,1]$. Thus, outside this range, the SILU function predominates.
Upon analyzing the splines learned by the network in different convolutions, we did not find any recognizable pattern. The behavior of the splines varies significantly across different convolutional layers and positions, indicating that the learning process is highly context-dependent and does not conform to a simple, uniform structure.
4.3 Parameters in a Convolutional KAN
As previously mentioned, the amount of parameters is one of the main advantages of using Convolutional KANs. With $\phi$ defined as in Equation 1, the parameters for each $\phi$ are the two weights, $w_{1}$ and $w_{2}$, together with the control points which can be adjusted to change the shape of each spline. Therefore there are $gridsize+2$ parameters for each $\phi$. Let the convolution kernel be of size $K\times K$, in total we have K²(gridsize + 2) parameters for each Convolutional KAN layer, compared to only K² for a CNN convolutional layer. In our experiments the gridsize is typically between k and k², with k tending to be a small value between 2 and 16.
In the convolutional layers, Convolutional KANs have more parameters, but as they utilize splines, they have more adaptability to process the spatial information and thus require less amount of fully connected layers which significantly increase the amount of parameters. That is where the advantage of utilizing splines really shows, we are able to reduce the amount of non-convolutional layers and thus reduce the parameter count.
5 Experiments
In this section we explain the different experiments we conducted to analyze the performance of different models that use KAN Convolutional Layers against a classical convolutional neural-network. During experimentation we used two datasets, MNIST and Fashion MNIST. We trained each architecture on both datasets to obtain the different models used to compare performance. For this we proposed architectures that use a mix between Linear, Kan Linear, Kan Convolutional and Convolutional layers. The following Figure 2 shows the different KAN Convolutional architectures used:
Figure 2: Convolutional KAN Architectures used in experiments
The following Figure 3 shows the different Standard architectures used in the experiments:
Figure 3: Standard Architectures used in experiments
Up to this version of the experiments, we have used the default splits of torch datasets for the training an test sets of the used datasets. Also, we have chosen hiperparameters and architectures by trial and error, but there is still place to conduct a thorugh hiperparameter search.
5.1 Loss
For every model that we trained, the Categorical Cross Entropy loss was used as base, but for KAN models, there are 2 additional regularization terms proposed in the KAN paper [6].
Cross entropy loss is defined as:
$L_{ce}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(p_{i,c})$ (6)
KAN loss with regularization is defined as:
$L_{reg}=L_{ce}+\lambda(\mu_{1}\sum_{l=1}^{L-1}|\Phi_{l}|_{1}+\mu_{2}\sum_{l=1}^{L-1}S(\Phi_{l}))$ (7)
With:
$S(\Phi_{l}))=-\sum_{i=1}^{n_{in}}\sum_{j=1}^{n_{out}}\frac{|\phi_{i,j}|}{|\Phi|_{1}}log(\frac{|\phi_{i,j}|}{|\Phi|_{1}})$ (8)
6 Results
This section displays an analysis of the performance of the different proposed models in the previously described experiments. The following Figure 4 shows the final accuracys of the proposed models.
Figure 4: Accuracy of the Architectures used in the MNIST (left) and Fashion MNIST (right) experiments. The proposed models all have better accuracy than CNN (Small) and are slightly worse than CNN (Medium) for both MNIST and Fashion MNIST
6.0.1 MNIST Dataset
Table 1 presents an analysis of performance metrics such as accuracy, precision, recall, F1 score, number of parameters, and training time per epoch for various models tested on the MNIST dataset. This analysis helps in understanding the efficiency and effectiveness of each model configuration.
Model Accuracy Precision Recall F1 Score #Params Minutes per epoch
MNIST Dataset
KKAN (Small) 98.90% 98.90% 98.89% 98.90% 94.875K 1.8119
Conv & KAN 98.75% 98.75% 98.75% 98.75% 95K 0.4551
KAN Conv & 1 Layer MLP 98.53% 98.54% 98.53% 98.53% 7.4K 1.7867
KAN Conv & 2 Layer MLP 98.58% 98.58% 98.58% 98.58% 164K 1.7779
KAN Conv BN & 2 Layer MLP 98.53% 98.54% 98.53% 98.53% 164K 1.7545
CNN (Medium) 99.12% 99.12% 99.12% 99.12% 157K 0.3110
CNN (Small) 97.59% 97.60% 97.59% 97.59% 2.7K 0.3056
1 Layer MLP 92.34% 92.22% 92.22% 92.22% 7.9K 0.2935
Table 1: Comparison of accuracy, precision, recall, F1 score, parameter count, and training time per epoch for the proposed models tested on the MNIST Dataset. Convolutional KANs altenative KKAN (Small) shows an increase in accuracy of $0.15\%$ against normal convolutions alternative with the same ammount of paramenters. Also KKAN with only $\sim 90$k parameters achieves $0.22\%$ less accuracy than the CNN (Medium) with 157k parameters.
Figure 5illustrates the relationship between test loss and parameter count versus accuracy for these models, providing a graphical representation of the data. The Figure 5 highlights how models with varying numbers of parameters perform in terms of test loss and accuracy, emphasizing the trade-offs involved in model complexity and performance.
Figure 5: Test loss against epochs (left) and Parameter count vs Accuracy (right) with MNIST Dataset
As seen in Table 1 and Figure 5, in the MNIST dataset, we find that the KKAN model with $\sim$ 90k parameters, achieves only $0.22\%$ less accuracy than the CNN (Medium), which has 157k parameters. While KAN Convolutions with a KAN Network seem to mantain performance while reducing the needed parameters, we find that using KAN Convolutions perform better than classic convolutions, going from $98.90\%$ accuracy vs $98.75\%$, showing that KAN Convolutions are effectively learning more.
If instead of using a KAN Network at the end we decide to use a MLP and vary the convolution type, results are still unclear as seen in the 2 following experiments:
1. In the comparison of KAN Conv & 1 layer MLP vs CNN(Small), where our model perfoms with 98.53% accuracy vs 97.59% the CNN.
2. In the comparison of KAN Conv & 2 layer MLP vs CNN(Medium), where our model perfoms with 98.58% accuracy vs 99.12% the CNN (Medium).
This shows that KANs seem to learn more in fewer parameters scenarios, but that is compensanted when addying more parameters to the MLP.
6.0.2 Fashion MNIST Dataset
As the Table 1 presented with the MNIST Dataset, the Table 2 presents a comparison of accuracy, precision, recall, F1 score, parameter count, and training time per epoch for the proposed models tested on the Fashion MNIST Dataset.
Model Accuracy Precision Recall F1 Score #Params Minutes per epoch
Fashion MNIST Dataset
KKAN (Small) 89.69% 89.62% 89.69% 89.62% 94.875K 1.8399
Conv & KAN 88.84% 88.76% 88.84% 88.77% 94.95K 0.3347
KAN Conv & 1 Layer MLP 88.57% 88.45% 88.57% 88.47% 7.385K 1.7765
KAN Conv & 2 Layer MLP 89.45% 89.46% 89.45% 89.45% 163.951K 1.7168
KAN Conv BN & 2 Layer MLP 89.08% 89.06% 89.08% 89.01% 163.786K 1.6535
CNN (Medium) 90.14% 90.13% 90.14% 90.13% 157.03K 0.2931
CNN (Small) 87.15% 86.99% 87.15% 87.00% 2.74K 0.2896
1 Layer MLP 84.33% 84.21% 84.33% 84.21% 7.85K 0.2778
Table 2: Comparison of accuracy, precision, recall, F1 score, parameter count, and training time per epoch for the proposed models tested on the Fashion MNIST Dataset. Accuracys in this Table show the same pattern as with with MNIST Dataset, showing that KKAN mantains a high accuracy of $89.69\%$ with $\sim 95$k parameters. In addition KKAN shows an improve in accuracy of $0.85\%$ against Conv & KAN, which only changes the KAN convolutions for classic convolutions.
Figure 6, illustrates the relationship between test loss and parameter count versus accuracy for these models, providing a graphical representation of the data.
Figure 6: On the left, we visualize the test loss vs the training epochs, on the right we have parameter count vs Accuracy in Fashion-MNIST dataset.
Following the results analysis, Figure 6 and Table 2 illustrate the relationship between parameter counts versus accuracy and test loss for the same models but applied to the Fashion MNIST dataset. In the experiments with this dataset, we find the same patterns as in MNIST, where using KAN alternatives result in better metrics.
In the case where we use a KAN Network after the convolutions, using KAN convolutions (KKAN Small) gets 89.69% accuracy while using classic convolutions (Conv & KAN model) gets $88.84\%$, but Conv & KAN is 6 times faster to train. When using a MLP at the end, KAN convolutions model KAN Conv & 1 layer MLP ($\sim 7k$ parameters) achieve $88.57\%$ accuracy, while CNN(Small) ($\sim 3$k parameters) only 87.15%.
When using a MLP as final block is used, we still get mixed results, where in the smaller case KAN Conv & 1 Layer gets $88.57\%$ accuracy, while CNN (Small) $87.15\%$. In the bigger case ($\sim 160$k parameters) KAN Conv & 2 Layer MLP gets $89.45\%$ and CNN (Medium) $90.14\%$.
Up to here, we compared alternatives with almost the same parameter count. We also find out that with much more less parameters, KAN alternatives show almost the same results. Comparing CNN(Medium) ($\sim 160k$ parameters) vs KKAN (Small) ($\sim 95$k parameters), we almost mantain accuracy, getting 0.5% less accuracy.
7 Conclusions
In this paper we proposed a new way to adapt the idea of learning splines proposed in Kolmogorov-Arnold Networks to convolutional layers used widely in Computer Vision. We implemented a KAN Convolutional Layer that uses a kernel made of learnable non-linear functions that use B-Splines. We have found out that we can almost mantain accuracy with fewer parameters than the models that use classical convolutions and showed empirically that KAN Convolutions are better learners. This is seen in the comparisons between the models that mantain the network after flattening the convolutional block output and only change the convolution type. Using a KAN Network at the end, KAN convolutions achieve higher accuracy. When trying an MLP, KAN convolutions achieve higher accuracy in the small models, but when having a 2 layer MLP, the classic CNN wins by 0.5% with $\sim 160$k parameters. But the key factor is that KANs seem to mantain accuracy with lower parameter count, reducing paramaters from $157$k to $95$k, the accuracy in Fashion Mnist is only reduced from $90.14\%$ to $89.69\%$, but the training time is almost 6 times slower with the current implementations of KANs and its derivatives.
Based on the original KAN paper [6] proposal that KANs are more interepretable, we have tried to find a way to interpret this new type of convolutions, but at the moment, we have not found any clear way to intepret the B-Splines learned in each pixel of the convolutions. Given that we are working on images, it seems that the classic approach of visualizing what the filter does to the image seems the more ’human’ way to get a sense of what is being learned.
The investigation leads us to understand the limitations of this new and promising idea. These limitations are similar to those presented in the original KAN paper. The KAN Linear Layer and our KAN Convolutional Layer are new and need to be optimized before they are able to be scaled properly, as shown in the time per epoch metrics. This paper is a starting point for integrating KANs into computer vision, and shows that Convolutional KANs have the potential to be an alternative to Convolutional Neural Networks.
8 Future work
This section describes the future experiments and optimizations we will be investigating. Most importantly, we will look to analyze the performance of all the models proposed with the Cifar10 and ImageNet datasets. This will allow us to properly analyze how KAN Convolutional Layers interpret more complex information. We will also be looking to implement updating the grid by increasing the number of control points. Interpretability is one of the strong points of KANs, as mentioned in the original KAN paper, we will investigate the resulting splines of the learning process and how to interpet this information in the context of Computer Vision. Finally we will be looking to optimize the KAN Convolutional and KAN Linear layers to be able to reduce the inference time and the time to train to be able to scale properly and test on large datasets.